OpenStudy (anonymous):

Given that the pair of triangles is similar, find the length of the side labeled x. Round to the nearest tenth, if necessary. 1st triangle is 33.5 & x 2nd triangle is 6.7 & 9.1

OpenStudy (anonymous):

33.5/x = 6.7/9.1 x = (33.5)(9.1)/6.7 x = 45.5 You are using basic proportions here